OpenStudy (anonymous):

Your friend just challenged you to a race through an obstacle course. You know in order to beat him, you must run 30 meters within 60 seconds in a southern direction. What does your average velocity need to be to win the race? (2 points) .5 meters per second, south 15 meters per second, south 2 meters per second, south 90 meters per second, south

OpenStudy (trojanpoem):

v = x/t = 30 / 70 = 0.5 meters per second

OpenStudy (anonymous):

@TrojanPoem your really smart.

OpenStudy (trojanpoem):

Thanks.